# Towards Foundation Models: Evaluation of Geoscience Artificial Intelligence with Uncertainty

### v1

@@ 7 Conclusion
We present an evaluation framework for current DLMs and future FMs that simultaneously considers performance uncertainty/variability, learning efficiency, and overlap between training and test data. We demonstrate our evaluation framework using PhaseNet, a popular phase picking DLM [47] to compare 3 training approaches (training from scratch and two transfer learning approaches). We implement an extensive training design comprising 5 models and 1,200 model instances over varying amounts of data from meticulously constructed data splits. We cast the results into a statistical framework which allows for direct comparison among models and training approaches across the training budgets.
We demonstrate the effectiveness and versatility of the evaluation framework by comparing classification and functional regression metrics. We show how no model is a global winner in terms of learning efficiency; that it depends on the total amount of training data used. We further demonstrate that training-test overlap must be considered in future FM evaluation as seen through performance improvement due to overlap. By providing estimates for the data and training uncertainty, a developer or practitioner can better target their efforts to mitigate sources they deem most important for their problem. We show that without the evaluation framework, there is a significant risk of model mis-ranking. Finally, our results demonstrate how transfer learning may not always be the best choice in limited data scenarios, which emphasizes the importance of continuing to compare FMs to their simpler DLM counterparts and inspect how training data budgets influence model performance. We hope that this research helps practitioners and developers by providing a general guide for what a truly revolutionary model might look like.
@@ 8 Acknowledgements
This work was performed under the US Department of Energy, National Nuclear Security Administration’s Office of Defense Nuclear Nonproliferation Research and Development. This work is approved for public release under LA-UR-24-31133. We would like to thank Yongsoo Park (Los Alamos National Laboratory), Lisa Linville (Sandia National Laboratory), and Leanna House (Virginia Tech) for their advice and/or edits.

### v2

@@ 8 Conclusion
We present an evaluation framework for seismic DLMs that simultaneously considers performance uncertainty and learning efficiency. We demonstrate our evaluation framework using PhaseNet, a popular phase picking DLM to compare 3 training approaches (training from scratch and 2 TL approaches). We implement an extensive training design utilizing deep ensembles and meticulously constructed training splits that partially controls for data leakage. We summarize the resulting 720 models instances through a statistical framework to compare the 3 training approaches across the training budgets with uncertainty.
We demonstrate the effectiveness and versatility of the evaluation framework by comparing classification and functional regression metrics. We show how no model is a global winner in terms of learning efficiency; that it depends on the total amount of training data used. By providing estimates for the data and training uncertainty, a developer or practitioner can better target their efforts to mitigate sources they deem most important for their problem. We show that without the evaluation framework, there is a significant risk of model mis-ranking. Finally, our results demonstrate how TL may not always be the best choice in limited data scenarios, which is particularly important as the seismic community invests in seismic FMs and emphasizes the importance to compare FMs to their simpler DLM counterparts and inspect how training data budgets influence model performance. We hope that this research helps practitioners and developers by providing an evaluation framework for robust model comparison.
@@ 9 Acknowledgements
This work was performed under the US Department of Energy, National Nuclear Security Administration’s Office of Defense Nuclear Nonproliferation Research and Development. This work is approved for public release under LA-UR-24-31133. We would like to thank Chengping Chai (Oak Ridge National Laboratory), Lisa Linville (Sandia National Laboratory), and Leanna House (Virginia Tech) for their advice and/or edits.
10 Appendix
10.1 Data Splits and Cluster Distributions
Figures A1, A2, A3, & A4, aim to provide a deeper view into how the clustering impacts the distributions of various signal properties of the waveforms. Since this work focuses primarily on P-picking, we use a 10 second window starting at the labeled P-arrival for these signal properties. In all figures, the testing set is shown in black, and each non-black curve coincides with the identically colored clusters in the main article (Figure 2). All curves are normalized for comparison such that they integrate to 1. A 10 second window was chosen because only about 60% of the 824,135 waveforms have a labeled S-wave, so an adaptive S-P interval window that captures only P-wave information was infeasible without further analysis. In addition, Figure A5 shows the distribution of S-P intervals for those waveforms that contain labeled S-arrivals. About 87% of the waveforms have S-P intervals less than 10 seconds, so most or all of the P-wave information is captured in the majority of the waveform, without including S-wave information for particularly long S-P intervals.
Figure A1: Distributions of the maximum log FFT amplitude within each 5 Hz frequency bin (rows) are shown for the 3 components (columns) for each cluster (non-black curves) and the testing set (black curve).
Figure A2: Distributions of all FFT amplitudes within each 5 Hz frequency bin (rows) are shown for the 3 components (columns) for each cluster (non-black curves) and the testing set (black curve).
Figure A3: Distributions of the frequencies coinciding with the maximum FFT amplitudes (i.e., the argmax) within a 10 second window starting at the labeled P-arrival are shown for the Z (left), N (middle), and E (right) components for each cluster (non-black curves) and the testing set (black curve).
Figure A4: Distributions of the log of the peak amplitudes within a 10 second window starting at the labeled P-arrival are shown for the Z (left), N (middle), and E (right) components for each cluster (non-black curves) and the testing set (black curve).
Figure A5: The distribution of the S-P intervals in seconds is shown for the INSTANCE data for which there are labeled S-arrivals. Only about 60% of the 824,135 waveforms contain a labeled S-arrival.